OpenStudy (anonymous):

What do the sin 25 and cosine 65 have in common. Explain why this is true.

OpenStudy (anonymous):

They are the same. Basically, sin and cos are complementary functions. and so are 25 and 65, 25 and 65 sum 90 °

OpenStudy (anonymous):

Refer to attached picture.

OpenStudy (anonymous):

Sry... Cos65 =sin25 Cos(90-25)=sin25

OpenStudy (anonymous):

The only tricky part is to recognize that because the right angle is 90, if one angle is 25, then the other has to be 65 to make the sum of the three equal to 180. Similarly, if one angle is 65, then the other has to be 25. What is sin(25) in that picture? Well, it's opposite over hypotenuse. A is opposite 25, so sin(25) = a/h What is cos(65) in that picture? Well, it's adjacent over hypotenuse. A is adjacent to 65, so cos(65) = a/h a/h = a/h, so sin(25) = cos(65)